OpenStudy (anonymous):

The amount of oil used by a ship traveling at a uniform speed varies jointly with the distance and the square of the speed. The ship uses 30 barrels of oil in traveling 85 miles at 42 mi/h. How many barrels of oil are used when the ship travels 26 miles at 54 mi/h? Round your answer to the nearest tenth of a barrel, if necessary.

OpenStudy (superdavesuper):

"The amount of oil used by a ship traveling at a uniform speed varies jointly with the distance and the square of the speed" so amount of oil = k*distance*speed^2 where k is a constant then u can use "30 barrels of oil in traveling 85 miles at 42 mi/h" to solve for k.....

OpenStudy (anonymous):

15.2?

OpenStudy (superdavesuper):

doesnt seem right - would u show me how u get 15.2 plz?

OpenStudy (anonymous):

30=k * 85*(42)^2 30 = 149940k k = 30/149940 = 3/14994 = 1/4998 d = 26 s = 54 O = (26•54^2)/4998 ≈ 15.17

OpenStudy (superdavesuper):

very good - i made a mistake in my rounding :( ur ans is right! so rounding to the nearest one tenth, it would be 15.2 barrel :)

OpenStudy (anonymous):

Thanks

OpenStudy (superdavesuper):

welcome n sorry for doubting ur ans at first!